This is from chapters of making sense of statistics 7th edition. End of chapter questions. There is no data. These are general questions.

  1. In a dependent Samples t test what is the df for a sample of 100 participants in a particular study?
  2. The dependent samples t test would have less sampling error in its design. Discuss the reasons.
  3. Given the results below what would you conclude about the mean difference between Group A and Group B.
  4. Give an example of one continuous data and one categorical data with two categories that would require performing a dependent sample t Test
Answer #1

The df for dependent sample t test is n-1= 100-1 = 99

The dependent sample t test has less sampling error because Paired t-test compares study subjects at 2 different times (paired observations of the same subject). Unpaired t-test (aka Student’s test) compares two different subjects. The paired t-test reduces intersubject variability (because it makes comparisons between the same subject), and thus is theoretically more powerful than the unpaired t-test.

Example for dependent sample t test with continuous data

After mnemonic training Before mnemonic training 204 223 393 412 391 402 265 285 326 353 220 243 423 342 443 340 480 582 464

For categorical data , the dependent sample test is conducted using McNrmar test . Example for this is

Drug 1 No Yes a 100 b 180 110 d 120 No 80 Drug 2 Yes 10 C 300 90 210